OpenStudy (anonymous):

x^2+5x-24=0 solve by completing the square use a comma to separate answers if necessary

OpenStudy (dumbcow):

Rewrite with constant on right side x^2 +5x = 24 Take middle coefficient, halve it then square it -->(5/2)^2 = 25/4 Add that to both sides x^2 +5x +25/4 = 24 + 25/4 Factor left side, notice it is now a perfect square (x + 5/2)^2 = 24 + 25/4 Combine right side (x + 5/2)^2 = 96/4 + 25/4 = 121/4 Take square root of both sides x + 5/2 = +-(11/2) Subtract 5/2 on both sides x = +-(11/2) - 5/2 Two solutions: x = 11/2 - 5/2 = 6/2 = 3 OR x = -11/2 - 5/2 = -16/2 = -8

hero (hero):

Dumb cow is a beast

OpenStudy (anonymous):

wow

hero (hero):

Would've looked much better in latex though

OpenStudy (dumbcow):

haha yeah didn't feel like it though panther hope that helped...same process can be used for any quadratic

OpenStudy (anonymous):

how about on my word problem i posted im really not getting it

OpenStudy (dumbcow):

The square root function is the inverse function of the square...so when you have a quadratic the way to undo the "x^2" to isolate x is to take the square root

OpenStudy (anonymous):

i think i can follow along but i get messed up all the time on the little differences in each problem i think i get one masted and then i get the curveball